Math 254A. Homework 5 Due Wednesday, 27 September

1. Exercise 3 on page 28. The solution should be nonzero.

2. Exercise 1 on page 34.

3. Exercise 5 on page 15. For full credit, reduce the number of different cases to consider (preferably just one).

4. Exercise 6 on page 38.

5. Exercise 7 on page 38.